# Monocular Open Vocabulary Occupancy Prediction for Indoor Scenes

### v1

@@ Abstract
Open-vocabulary 3D occupancy is vital for embodied agents, which need to understand complex indoor environments where semantic categories are abundant and evolve beyond fixed taxonomies. While recent work has explored open-vocabulary occupancy in outdoor driving scenarios, such methods transfer poorly indoors, where geometry is denser, layouts are more intricate, and semantics are far more fine-grained. To address these challenges, we adopt a geometry-only supervision paradigm that uses only binary occupancy labels (occupied vs. free). Our framework builds upon 3D Language-Embedded Gaussians, which serve as a unified intermediate representation coupling fine-grained 3D geometry with a language-aligned semantic embedding. On the geometry side, we find that existing Gaussian-to-Occupancy operators fail to converge under such weak supervision, and we introduce an opacity-aware, Poisson-based approach that stabilizes volumetric aggregation. On the semantic side, direct alignment between rendered features and open-vocabulary segmentation features suffers from feature mixing; we therefore propose a Progressive Temperature Decay schedule that gradually sharpens opacities during splatting, strengthening Gaussian–language alignment. On Occ-ScanNet, our framework achieves 59.50 IoU and 21.05 mIoU in the open-vocabulary setting, surpassing all existing occupancy methods in IoU and outperforming prior open-vocabulary approaches by a large margin in mIoU. Code will be released at https://github.com/JuIvyy/LegoOcc.
@@ 1 Introduction
To address this gap, we adopt a geometry-only supervision protocol: training uses only binary voxel occupancy labels and no semantic voxel annotations, and we tackle the resulting challenges on both the geometric and semantic fronts. This setting is practically motivated, as geometry can be harvested at scale with modest effort—for example, Occ-ScanNet constructs its occupancy ground truth via SCFusion [46], which incrementally fuses occupancy probabilities from depth-based reconstruction [40] and completion methods [3, 43]. In contrast, semantic labels remain expensive to curate due to the high diversity and long-tailed distribution of indoor categories.
@@ 2.2 Open Vocabulary Occupancy Prediction
Open-vocabulary occupancy has recently gained traction in autonomous driving as a way to mitigate the cost and difficulty of annotating 3D occupancy. These methods typically adopt self-/weakly supervised settings to learn voxel-level geometry and semantics without full 3D labels. For semantic learning, most approaches first derive open-vocabulary cues from VLMs and obtain pseudo 2D masks with open-set segmentors [31, 24]. These signals are then either lifted to 3D by projecting pixels onto point clouds or associating pixels/points with voxels [54, 19, 53], or mapped to voxels via differentiable voxel projection [58, 41], or used to supervise a rendered 3D semantic field directly in the image plane via volume rendering [5]. Feature alignment is commonly achieved through cosine similarity between voxel features and text/image embeddings [31, 24]. For geometry learning, two families dominate: LiDAR-derived pseudo labels aggregated into voxels for supervision [58, 9, 41, 19], and self-supervised multi-view rendering to learn occupancy by enforcing photometric/feature consistency across frames without 3D ground truth [11, 54].
@@ 3.1 Problem Setting
Figure 2: LegoOcc Framework Overview. From a monocular image, a feed-forward Gaussian model produces Language-Embedded Gaussians. Training proceeds along two couched paths: Semantic learning, we differentiably render Gaussian features to the image with Progressive Temperature Decay and align them to a training-free open-vocabulary segmenter via a cosine objective $L_{\text{feat}}$; Geometry learning, we convert Gaussians to occupancy using a opacity-aware Poisson-based Gaussian-to-Occupancy operator and supervise binary occupancy with $L_{\text{occ}}$. At inference, the Language-Embedded Occupancy supports text-driven queries by computing cosine similarity between voxel embeddings and prompt embeddings, yielding open-vocabulary semantic occupancy without dense voxel-level semantic labels during training.
@@ 3.3 Poisson-based Gaussian-to-Occupancy
Gaussian Splatting. We follow the front-to-back $\alpha$-blending strategy of Gaussian Splatting [17] to render image features from LE-Gaussians. For the $i$-th Gaussian $\mathcal{G}_{i}$ with color $\mathbf{c}_{i}$, opacity $\alpha_{i}\!\in[0,1]$, and screen-space mean/covariance $(\bm{\mu}^{\prime}_{i},\bm{\Sigma}^{\prime}_{i})$, its 2D kernel at image location $\mathbf{x}^{\prime}$ is
@@
While efficient and conceptually simple, this formulation omits opacity $\alpha_{i}$ in the geometry branch and relies solely on the spatial kernels. As a result, the occupancy inference is decoupled from the opacity used in image rendering, creating a mismatch between 2D rendering and 3D voxel aggregation.
@@
However, we observe that the Bernoulli aggregation drives the learned opacities to small values, which enlarges the gap between rendered features and per-Gaussian embeddings (see Sec. 3.4). This behavior likely arises because when multiple Gaussians overlap within a voxel, the Bernoulli union $1-\prod_{i}(1-\alpha_{i}p_{i})$ saturates rapidly toward 1; and for a fixed total mass $\sum_{i}\alpha_{i}p_{i}$, the union is maximized when the each of the overlapping contributions are similar, making it difficult to learn differentiated opacities.
@@ 3.4 Progressive Temperature Decay
As shown in the semantic learning part of Fig. 2, we adopt Gaussian Splatting [17] to render LE-Gaussians into images. For semantic learning, we render LE-Gaussian features via Gaussian Splatting [17] to images. However, the $\alpha$-blending in Eq. 4 produces a weighted sum of per-ray contributions, so each pixel feature becomes a mixture of multiple Gaussian embeddings along the ray. Consequently, the supervision signal encourages the rendered mixture, rather than each individual Gaussian, to be language-aligned, creating a mismatch that later hampers G2O-based splatting. This ambiguity is further exacerbated in indoor scenes, where objects are numerous, fine-grained, and heavily overlapping in projection. Dr. Splat [16] mitigates this via hard feature registration: along a pixel ray it assigns the pixel’s CLIP embedding to only the Top-$k$ dominant Gaussians, yielding language-aligned vectors that are directly retrievable in 3D. While effective, this discrete selection causes sparse gradients, since supervision backpropagates only through the chosen Top-$k$ primitives.
Instead of a hard Top-$k$, we reduce mixture ambiguity via a continuous and training-stable opacity sharpening. Let $\alpha_{i}=\sigma(\alpha_{i}^{\text{logit}})$, where $\sigma$ is the sigmoid function and $\alpha_{i}^{\text{logit}}$ is the model-predicted logit. We adopt a tempered sigmoid
@@
where smaller (near 0) temperature $\tau$ values sharpen opacities towards $\{0,1\}$, reducing feature mixing along the ray while maintaining gradient flow to all contributors (rather than a discrete subset). In practice, we employ a progressive temperature decay schedule during training, annealing $\tau$ from $T_{\max}$ to $T_{\min}$ (default $T_{\max}{=}1$, $T_{\min}{=}10^{-3}$). Let $r\!\in[0,1]$ denote the normalized training progress, 0 means begnning, 1 means end. We use an exponential schedule
@@
Table 1: Monocular results on Occ-ScanNet. We re-implement POP-3D [41] and LOcc [53] for the indoor monocular setting, and supervise geometry with ground-truth binary occupancy instead of the authors’ LiDAR-derived pseudo labels. For POP-3D, we project voxel centers to the image plane and sample language-aligned features at those locations to form a 3D grid of text-aligned embeddings for open-vocabulary reasoning. For LOcc, we prompt Qwen2.5-VL-7B [4] to extract object names and use Trident [35] to obtain training-free open-vocabulary segmentations, which yield stronger 2D masks on Occ-ScanNet. Adopt DepthAnything-v2 [48] as the backbone and use the same input resolution and training recipe as ours for fair comparison.
@@ 3.5 Losses
We optimize the network with a composite objective that couples 3D geometry supervision with 2D language-aligned guidance as shown in Fig. 2. For occupancy, we follow [47] but use binary voxel supervision with focal loss $L_{\mathrm{focal}}$ and Lovász–Softmax $L_{\mathrm{lov}}$, plus a scene-class affinity regularizer $L_{\mathrm{scal}}$ for spatial coherence. Geometry is further stabilized by a robust Huber depth loss $L_{\mathrm{depth}}$. For semantics, we align image features from LE-Gaussians with with frozen open-vocabulary segmentation features via a cosine objective $L_{\mathrm{feat}}$. To exploit multi-view consistency without 2D labels, we re-render nearby views (default: 5 frames) and apply the same feature alignment. The final objective is
@@ 4.2 Implementation Details
We adopt Depth-Anything V2 [48] as our depth backbone, which has been used as a geometry prior in [50, 47]. Building on this, we construct Gaussian primitives using a surface-point expansion strategy similar to [2]. All models are trained for 10 epochs with a batch size of 4, optimizing with AdamW [26]; the learning rate is initialized at $2\times 10^{-4}$ and follows a cosine decay schedule [25] with linear warmup, with weight decay set to $0.01$. We apply gradient clipping [27] with a maximum norm of $1.0$, and resize each image so that its longer side is $518$ pixels while preserving aspect ratio. We train all our model using 4 NVIDIA GeForce RTX 4090 GPUs.
@@ 4.3 Main Results
From Tab. 1, we observe that using geometry-only annotation the re-implemented baselines POP-3D [41] and LOcc [53] perform poorly, obtains only 35.32/5.96 and 36.70/9.25 IoU/mIoU, respectively. In contrast, our model reaches 59.50 IoU and 21.05 mIoU, outperforming both by a large margin. Notably, LegoOcc achieves the highest IoU among all entries, indicating that the proposed design effectively recovers occupancy without semantic voxel labels. The remaining performance gap between full-annotation methods and geometry-only ones in mIoU highlights the intrinsic difficulty of category calibration for open-vocabulary occupancy, particularly under class ambiguity. We further illustrate these problem in our appendix.
@@ 4.4 Ablation Studies
Model mIoU IoU FPS
ISO [50] 28.71 42.16 3.81
EmbodiedOcc [47] 45.15 53.55 11.48
POP-3D [41] 5.96 35.32 10.21
LOcc [53] 9.25 36.70 8.93
Ours 21.05 59.50 22.47

### v2

@@ Abstract
Open-vocabulary 3D occupancy is vital for embodied agents, which need to understand complex indoor environments where semantic categories are abundant and evolve beyond fixed taxonomies. While recent work has explored open-vocabulary occupancy in outdoor driving scenarios, such methods transfer poorly indoors, where geometry is denser, layouts are more intricate, and semantics are far more fine-grained. To address these challenges, we adopt a geometry-only supervision paradigm that uses only binary occupancy labels (occupied vs. free). Our framework builds upon 3D Language-Embedded Gaussians, which serve as a unified intermediate representation coupling fine-grained 3D geometry with a language-aligned semantic embedding. On the geometry side, we find that existing Gaussian-to-Occupancy operators fail to converge under such weak supervision, and we introduce an opacity-aware, Poisson-based approach that stabilizes volumetric aggregation. On the semantic side, direct alignment between rendered features and open-vocabulary segmentation features suffers from feature mixing; we therefore propose a Progressive Temperature Decay schedule that gradually sharpens opacities during splatting, strengthening Gaussian-language alignment. On Occ-ScanNet, our framework achieves 59.50 IoU and 21.05 mIoU in the open-vocabulary setting, surpassing all existing occupancy methods in IoU and outperforming prior open-vocabulary approaches by a large margin in mIoU. Code will be released at https://github.com/JuIvyy/LegoOcc.
@@ 1 Introduction
To address this gap, we adopt a geometry-only supervision protocol: training uses only binary voxel occupancy labels and no semantic voxel annotations, and we tackle the resulting challenges on both the geometric and semantic fronts. This setting is practically motivated, as geometry can be harvested at scale with modest effort. Specifically, ScanNet provides images and depth captured by an RGB-D camera, and Occ-ScanNet generates occupancy ground truth via the SCFusion pipeline [46], which incrementally fuses occupancy probabilities from depth-based reconstruction [40] and completion methods [1, 43]. In particular, SCFusion fuses depth with aligned ShapeNet CAD models via Scan2CAD [1] to recover complete-scene geometry and mitigate occlusions inherent in raw sensor data. In contrast, semantic labels remain expensive to curate due to the high diversity and long-tailed distribution of indoor categories.
@@ 2.2 Open Vocabulary Occupancy Prediction
Open-vocabulary occupancy has recently gained traction in autonomous driving as a way to mitigate the cost and difficulty of annotating 3D occupancy. These methods typically adopt self-/weakly supervised settings to learn voxel-level geometry and semantics without full 3D labels. For semantic learning, most approaches first derive open-vocabulary cues from VLMs and obtain pseudo 2D masks with open-set segmentors [31, 23, 13]. These signals are then either lifted to 3D by projecting pixels onto point clouds or associating pixels/points with voxels [54, 18, 53], or mapped to voxels via differentiable voxel projection [58, 41], or used to supervise a rendered 3D semantic field directly in the image plane via volume rendering [3]. Feature alignment is commonly achieved through cosine similarity between voxel features and text/image embeddings [31, 23]. For geometry learning, two families dominate: LiDAR-derived pseudo labels aggregated into voxels for supervision [58, 7, 41, 18], and self-supervised multi-view rendering to learn occupancy by enforcing photometric/feature consistency across frames without 3D ground truth [9, 54].
@@ 3.1 Problem Setting
Figure 2: LegoOcc Framework Overview. From a monocular image, a feed-forward Gaussian model produces Language-Embedded Gaussians. Training proceeds along two paths: Semantic learning, we differentiably render Gaussian features to the image with Progressive Temperature Decay and align them to a training-free open-vocabulary segmenter via a cosine objective $L_{\text{feat}}$; Geometry learning, we convert Gaussians to occupancy using an opacity-aware Poisson-based Gaussian-to-Occupancy operator and supervise binary occupancy with $L_{\text{occ}}$. At inference, the Language-Embedded Occupancy supports text-driven queries by computing cosine similarity between voxel embeddings and prompt embeddings, yielding open-vocabulary semantic occupancy without dense voxel-level semantic labels during training.
@@ 3.3 Poisson-based Gaussian-to-Occupancy
Gaussian Splatting. We follow the front-to-back $\alpha$-blending strategy of Gaussian Splatting [16] to render features from LE-Gaussians. For the $i$-th Gaussian $\mathcal{G}_{i}$ with feature $\mathbf{f}_{i}$, opacity $\alpha_{i}$, and screen-space mean/covariance $(\bm{\mu}^{\prime}_{i},\bm{\Sigma}^{\prime}_{i})$, its 2D kernel at image location $\mathbf{x}^{\prime}$ is
@@
While efficient and simple, this formulation omits opacity $\alpha_{i}$ in the geometry branch and relies solely on the spatial kernels. As a result, the occupancy inference is decoupled from the opacity used in image rendering, creating a mismatch between 2D rendering and 3D voxel aggregation.
@@
However, we observe that the Bernoulli aggregation drives the learned opacities to small values, which enlarges the gap between rendered features and per-Gaussian embeddings (see Sec. 3.4). This behavior likely arises because when multiple Gaussians overlap within a voxel, the Bernoulli union $1-\prod_{i}(1-\alpha_{i}p_{i})$ saturates rapidly toward 1; and for a fixed total mass $\sum_{i}\alpha_{i}p_{i}$, the union is maximized when each of the overlapping contributions are similar, making it difficult to learn differentiated opacities.
@@ 3.4 Progressive Temperature Decay
As shown in the semantic learning part of Fig. 2, we adopt Gaussian Splatting [16] to render LE-Gaussians into images. However, the $\alpha$-blending in Eq. 4 produces a weighted sum of per-ray contributions, so each pixel feature becomes a mixture of multiple Gaussian embeddings along the ray. Consequently, the supervision signal encourages the rendered mixture, rather than each individual Gaussian, to be language-aligned, creating a mismatch that later hampers G2O-based splatting. This ambiguity is further exacerbated in indoor scenes, where objects are numerous, fine-grained, and heavily overlapping in projection. Dr. Splat [15] mitigates this via hard feature registration: along a pixel ray it assigns the pixel’s CLIP embedding to only the Top-$k$ dominant Gaussians, yielding language-aligned vectors that are directly retrievable in 3D. While effective, this discrete selection causes sparse gradients, since supervision backpropagates only through the chosen Top-$k$ primitives.
Instead of a hard Top-$k$, we reduce mixture ambiguity via a continuous and training-stable opacity sharpening. Let $\alpha_{i}=\sigma(\alpha_{i}^{\text{logit}})$, where $\sigma$ is the sigmoid function and $\alpha_{i}^{\text{logit}}$ is the predicted logit. We adopt a tempered sigmoid
@@
where smaller (near 0) temperature $\tau$ values sharpen opacities towards $\{0,1\}$, reducing feature mixing along the ray while maintaining gradient flow to all contributors (rather than a discrete subset). In practice, we employ a progressive temperature decay schedule during training, annealing $\tau$ from $T_{\max}$ to $T_{\min}$ (default $T_{\max}{=}1$, $T_{\min}{=}10^{-3}$). Let $r\!\in[0,1]$ denote the normalized training progress, 0 means beginning, 1 means end. We use an exponential schedule
@@
Table 1: Monocular results on Occ-ScanNet. We re-implement POP-3D [41] and LOcc [53] for the indoor monocular setting, and supervise geometry with ground-truth binary occupancy instead of the authors’ LiDAR-derived pseudo labels. We also adopt DepthAnything-v2 [48] as the backbone and use the same input resolution and training recipe as ours for fair comparison. For POP-3D, we project voxel centers to the image plane and sample language-aligned features at those locations to form a 3D grid of text-aligned embeddings for open-vocabulary reasoning. For LOcc, we prompt Qwen2.5-VL-7B [2] to extract object names and use Trident [35] to obtain training-free open-vocabulary segmentations, which yield stronger 2D masks on Occ-ScanNet.
@@ 3.5 Losses
We optimize the network with a composite objective that couples 3D geometry supervision with 2D language-aligned guidance as shown in Fig. 2. For occupancy, we follow [47] but use binary voxel supervision with focal loss $L_{\mathrm{focal}}$ and Lovász–Softmax $L_{\mathrm{lov}}$, plus a scene-class affinity regularizer $L_{\mathrm{scal}}$ for spatial coherence. Geometry is further stabilized by a robust Huber depth loss $L_{\mathrm{depth}}$. For semantics, we align image features from LE-Gaussians with frozen open-vocabulary segmentation features via a cosine objective $L_{\mathrm{feat}}$. To exploit multi-view consistency without 2D labels, we re-render nearby views (default: 5 frames) and apply the same feature alignment. The final objective is
@@ 4.2 Implementation Details
We adopt Depth-Anything V2 [48] as our depth backbone, which has been used as a geometry prior in [50, 47]. Building on this, we construct Gaussian primitives using a surface-point expansion strategy similar to [59]. For semantics, we use Trident [35] as the open-vocabulary segmentation model to extract language-aligned per-pixel features, and the CLIP [30] text encoder to embed category names. We follow the standard prompt template “a photo of $\{\cdot\}$” to obtain text embeddings.
All models are trained for 10 epochs with a batch size of 4, optimizing with AdamW [25]; the learning rate is initialized at $2\times 10^{-4}$ and follows a cosine decay schedule [24] with linear warmup, with weight decay set to $0.01$. We apply gradient clipping [27] with a maximum norm of $1.0$, and resize each image so that its longer side is $518$ pixels while preserving aspect ratio. We train all our model using 4 NVIDIA GeForce RTX 4090 GPUs.
@@ 4.3 Main Results
From Tab. 1, we observe that using geometry-only annotation the re-implemented baselines POP-3D [41] and LOcc [53] perform poorly, obtaining only 35.32/5.96 and 36.70/9.25 IoU/mIoU, respectively. In contrast, our model reaches 59.50 IoU and 21.05 mIoU, outperforming both by a large margin. Notably, LegoOcc achieves the highest IoU among all entries, indicating that the proposed design effectively recovers occupancy without semantic voxel labels. The remaining performance gap between full-annotation methods and geometry-only ones in mIoU highlights the intrinsic difficulty of category calibration for open-vocabulary occupancy, particularly under class ambiguity. We further illustrate these problems in our appendix.
@@ 4.4 Ablation Studies
Model mIoU IoU FPS #parameters
ISO [50] 28.71 42.16 3.81 303.05M
EmbodiedOcc [47] 45.15 53.55 11.48 231.45M
POP-3D [41] 5.96 35.32 10.21 115.31M
LOcc [53] 9.25 36.70 8.93 113.50M
Ours 21.05 59.50 22.47 99.04M
@@ 5 Conclusion
Acknowledgements
This work was supported by National Natural Science Foundation of China (NFSC) under the Grant Number 62573370 and Key Area Project of Education Department of Guangdong Province (No. 2025ZDZX3051).